# Customize-A-Video: One-Shot Motion Customization of Text-to-Video Diffusion Models

### v1

@@ C.3 Future Plans
Abundant spatial customization approaches have been developed for T2I diffusion models. We leverage some of them to serve as our appearance absorbers for their training stability on few-shot learning and inference simplicity in the staged scheme. In the next step we plan to investigate more options to discover their characteristics and further enhance our method’s performance and usability. Besides, video diffusion models are also rapidly evolving and our modules are inherently compatible with various types of temporal attentions regardless of the input modalities including image-to-video and video-to-video tasks.

### v3

@@ Appendix 0.E Future Work
Abundant image customization approaches with various tuning techniques have been developed for T2I diffusion models. We leverage some of them to serve as our appearance absorbers for their training stability on few-shot learning and inference simplicity in the staged scheme. In the next step we plan to investigate more options to discover their characteristics and further enhance our method’s performance and usability. Besides, generative video foundation models are also rapidly evolving and our modules are inherently compatible with various types of temporal attentions, regardless of the specific generation process and input modalities.